OpenStudy (anonymous):

At the city museum, child admission is $6.40 and adult admission is $9.50. On Thursday, 151 tickets were sold for a total sales of $1239.20. How many adult tickets were sold that day?

OpenStudy (anonymous):

Let's have x be the number of children tickets and 2x will be the adults. Now we can set up the following equation: 5.4x+8.9(2x)=626.4 Expand this out: 5.4x+17.8x=626.4 Combine like terms: 23.2x=626.4 Divide both sides by 23.2 to get: x=27 There we go, 27 children tickets were sold.

OpenStudy (sushi121212):

The answer is 88. You have to set the tickets equal to each other: a=adults c=children 6.40c+9.50a=1239.20 c+a=151 Then, you have to use substitution or elimination: -6.40 x c+a=151 ( everything gets multiplied into everything ) 6.40c+9.50a=1239.20 -6.40c-6.40a=966.4 ______________________ (subtract) 3.10a = 272.8 _____ ______ 3.10 3.10 a=88 That means that there were 88 adult tickets purchased.

OpenStudy (anonymous):

LOL You got the adult tickets and i got the children tickets

OpenStudy (sushi121212):

Haha, that's funny! What a coincidence!